OpenStudy (anonymous):

The county animal park is shaped like a triangle. Two perpendicular sides of the park measure 2.3 miles and 3.4 miles. What is the area of the animal park? A. 3.91 mi² B. 5.7 mi² C. 7.82 mi² D. 11.4 mi²

OpenStudy (anonymous):

if you give me answer I will give you medal im new to open study

OpenStudy (anonymous):

A. 3.91 mi^2

OpenStudy (anonymous):

Surface of a triangle is base*height/2

OpenStudy (anonymous):

ok hope ur right

OpenStudy (anonymous):

say base = 2.3 mi and height = 3.4 mi, then S = 2.3*3.4/2 = 3.91 mi^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im on my last question

OpenStudy (anonymous):

What is the area of a triangle with a base of 2.4 cm and a height of 0.8 cm? A. 0.96 cm² B. 1.6 cm² C. 1.92 cm² D. 3.2 cm²

OpenStudy (anonymous):

can you tell me the answer to this one to lol XD

OpenStudy (anonymous):

use the formula i just told you, area = base*height/2 = 2.4*0.8/2 = 0.96 cm^2 (A)

OpenStudy (anonymous):

right sorry